OpenStudy (anonymous):

Find T, N and kappa r(t) = (2t+3)i +(5-t^2)j

zepdrix (zepdrix):

Ooo I haven't done these in a while :U a bit rusty. So ummm r(t) is a position vector. So the tangent vector is the derivative of positionnnnnnnnn? And then to get the `unit` tangent vector, we divide by it's magnitude. \[\Large\bf\sf \hat T\quad=\quad \frac{\vec r\;'}{|\vec r\;'|}\]Does that look right, maybe?

OpenStudy (anonymous):

I am only having problems with one part of this problem. that will make it shorter to help

zepdrix (zepdrix):

Oh which part XD

OpenStudy (anonymous):

when finding T, specifically the j component.

OpenStudy (anonymous):

im sorry, not T.. for finding N

OpenStudy (anonymous):

specifically dT/dt / |dT/dt| for the j component.

OpenStudy (anonymous):

and the problem is taking the derivative of T for the j component. I am stuck on the simplification. I can leave it in a more expanded form.. but that is really nasty to deal with when dividing by the length to get N.

zepdrix (zepdrix):

Let's make sure we have the same Tangent Vector first of all. Did you come up with something like this?\[\Large\bf\sf \hat T\quad=\quad \left<\frac{1}{\sqrt{1+t^2}},\;\frac{-t}{\sqrt{1+t^2}}\right>\]

OpenStudy (anonymous):

exactly

zepdrix (zepdrix):

Oh boy, quotient rule huh? :( this is gonna be a doozy.

zepdrix (zepdrix):

I went through this kinda quick, maybe made a mistake,\[\Large\bf\sf \hat T'\quad=\quad \left<\frac{-t}{(1+t^2)^{3/2}},\;\frac{-1}{(1+t^2)^{3/2}}\right>\]Does this look rightttt?

OpenStudy (anonymous):

i will show you where I get to and where I am trying to get .\[\frac{ \sqrt{1+t^2}-t(\frac{ t }{ \sqrt{1=t^2} }) }{ (\sqrt{1+t^2})^3 }\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

how do get the above into your j component.

zepdrix (zepdrix):

How do you what? :U

zepdrix (zepdrix):

You mean, how do you simplify after quotient rule?

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

Yes I think that's the correct simplification :o Need help getting to that?

OpenStudy (anonymous):

yes!! :)

zepdrix (zepdrix):

From here, \[\Large\bf\sf \frac{-\sqrt{1+t^2}-\left(\frac{-t^2}{\sqrt{1+t^2}}\right)}{(1+t^2)}\]Get a common denominator up top.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{-\dfrac{(1+t^2)}{\sqrt{1+t^2}}-\left(\dfrac{-t^2}{\sqrt{1+t^2}}\right)}{(1+t^2)}\]

OpenStudy (anonymous):

is there supposed to be a root in the denominator?

zepdrix (zepdrix):

I multiplied the first term by sqrt(1+t^2), top and bottom, to get a common denominator with the other fraction.

OpenStudy (anonymous):

i see it now!!!

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

Awesome.. a shade under 2 hours for this one :) well just to figure out that part.. the rest was easy enough.

OpenStudy (anonymous):

thanks for your help tremendously!!

zepdrix (zepdrix):

ok cool \c:/ good job